What is x in the given surface area of a right prism?
S=320 ft2
S=2(lw+hl+hw)
lw= x
hl= 8 ft
hw= 6 ft

Answer by Alan3354(69443)

What is x in the given surface area of a right prism?
S=320 ft2
S=2(lw+hl+hw)
lw= x
hl= 8 ft
hw= 6 ft
----------------
S=2(lw+hl+hw) = 320
(lw+hl+hw) = 160
x + 8 + 6 = 160
x = 144 ft